Suppose GDP is Rs.10.3 trillion, taxes are Rs.1.8 trillion, private saving is Rs.1.2 trillion, and public saving is Rs.0.2 trillion. Assuming this economy is closed, calculate


1
Expert's answer
2021-06-03T13:12:42-0400

Private savings= GDP-taxes-consumption


1.2= 10.3"-" 1.8-consumption


Consumption= 10.3"-" 1.8"-" 1.2 = 7.3


National savings = public + private savings


= 1.2+0.2 = 1.4


NOW,


National savings = GDP"-" Consumption- govt purchases


1.4 = 10.3 "-" 7.3 =govt purchases


Govt purchases = 10.3"-" 7.3"-" 1.4 = 1.6


Now we know


GDP = C + I + G


10.3 = 7.3 + I + 1.6


I = 10.3-7.3"-" 1.6 = 1.4